A 3.00L container contains 326.14 g of chlorine gas, 2.50 mol of iodine gas and 1.92 mol of neon gas. The total pressure inside the container was measured to be 6.78 atm. What is the partial pressure of chlorine gas?
1
Expert's answer
2020-03-24T14:01:25-0400

The number of moles of chlorine gas in the contaner can be calculated as following:

n = m/Mr,

where n - number of moles, m - mass, Mr - molecular weight.

From here:

n(Cl2) = 326.14 g / 71 g/mol = 4.6 mol

Partial pressure of chlorine gas is proportional to the molar fraction of chlorine gas in the container:

P(Cl2) = X(Cl2) × P(total)

where P(Cl2) - partial pressure of chlorine gas, X(Cl2) - molar fraction of chlorine gas, P(total) - total pressure.

X(Cl2) = n(Cl2) / [n(Cl2) + n(I2) + n(Ne)] = 4.6 mol / [4.6 mol + 2.5 mol + 1.92 mol] = 0.51 mol

As a result:

P(Cl2) = 0.51 mol × 6.78 atm = 3.46 atm


Answer: 3.46 atm